OpenStudy (anonymous):

imaginary numbers... can someone tell me if im correct. im not sure if im understanding this yet. so if the power of "i" is divisible by 4, the answer would be a positive 1 Ex: i^4 , i^80 = 1 , i^24= 1 if the power of "i" is divisible by 2 but not by 4, the answer would be a negative 1 Ex: i^22 , i^2 , i^14 if the power of "i" is an odd number then the answer would be a negative "i" Ex; i^3 , i^27 , i^67 & finally, where would positive "i" fit in?

OpenStudy (anonymous):

yep. what do you mean by "positive i"?

OpenStudy (anonymous):

yes you have it exactly look only at the remainder and match up with \(i^1=i,i^2=-1,i^3=-i,i^4=1\)

OpenStudy (anonymous):

however there is no such thing as "positive \(i\) because complex numbers are neither positive or negative. it is just "\(i\)" or "minus \(i\)

OpenStudy (anonymous):

oh well yeah tht is what i ment, i

OpenStudy (anonymous):

so if the power of i is a negative number then the answer would be i ?

OpenStudy (anonymous):

a negative power is equivalent to saying 1 divided by the value raised to "opposite" power (i.e., 1/(x^3) for x^(-3)

OpenStudy (anonymous):

okay so can u give me examples that their answers would be "i" ? i think that would help me. please.

OpenStudy (anonymous):

"negative exponent" does not imply that there is an "i" in the expression

OpenStudy (anonymous):

yes i understand that.

OpenStudy (anonymous):

so how would you solve this